OpenStudy (anonymous):

I NEED MATH HOMEWORK HELP I HAVE 10 QUESTIONS I NEED ANSWERED PLEASE HELP!!!!! I DON'T UNDERSTAND IT!!!!!!!!!!!!!!!

OpenStudy (blurbendy):

What is/are your question(s)?

OpenStudy (anonymous):

Simplify the expression. 2 exponent 4 A. -16 B -8 C 8 D 16

OpenStudy (anonymous):

2 x 2 x 2 x 2 = 16 :)

OpenStudy (anonymous):

Thanks! Could you help me with another one?

OpenStudy (anonymous):

Yes sure

OpenStudy (anonymous):

Simplify the expression. 12 exponent 0 0 1 12 120

OpenStudy (anonymous):

Whatever number with the exponent 0 is ALWAYS 1. Remember that as a tip!

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

Youre welcome :D any more?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Write as a single power of the given base. 4 exponent 2 + 3 exponent 3 17 35 43 97

OpenStudy (raden):

except, 0^0 cant be equals 1 @kailefinies ?

OpenStudy (anonymous):

It is

OpenStudy (anonymous):

4^2 + 3^3 = (4 x 4) + (3 x 3 x 3) = 16 + 27 = 43

OpenStudy (anonymous):

Thanks so much! Evaluate the expression if a = 2. 5a exponent 4 30 40 60 80

OpenStudy (anonymous):

\[5a^4\] you then fill in a = 2 \[5 * 2^4\] that becomes \[5 * 16= 80\]

OpenStudy (anonymous):

You're a big help! Thanks so much! this is my last question. Which is 625 written as a power of a number? 3 exponent 4 3 exponent 5 5 exponent 3 5 exponent 4

OpenStudy (anonymous):

Alright, try writing them all out on a piece of paper. \[3^4 = 3*3*3*3\] \[3^5= 3*3*3*3*3\] \[5^3 = 5*5*5\] \[5^4= 5*5*5*5\] then look at your answers and see which one becomes 625, goodluck!

OpenStudy (anonymous):

Thank you so much!!!! Yo have been a HUGE help! I can't thank you enough!

OpenStudy (anonymous):

Youre welcome :) I'm glad I could help!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Oh no! I ran into more problems!

OpenStudy (anonymous):

Can you help please????

OpenStudy (anonymous):

Yessss, sorry for the late response, but im here now

OpenStudy (anonymous):

What is the prime factorization of 54? A 9 • 6 B 2 exponent 2 • 3 exponent 2 C 3 exponent 3• 2 D 3 exponent 2• 6

OpenStudy (anonymous):

The answer is C) \[3^3 * 2\] (3 *3 * 3) * 2 = 27 * 2 = 54 For future problems like these just grab your calculator and calculate the answers to see which one matches ;)

OpenStudy (anonymous):

Thank you thank you thank you!! Another question, How do I calculate exponent problems like this on a calculator????

OpenStudy (anonymous):

do you have a normal calculator or a graphic one?

OpenStudy (anonymous):

I have a TI-34

OpenStudy (anonymous):

alright, the first option that will work on any calculator is: for example you need to know 3 exponent 4, you type in your calculator: 3 * 3 * 3 * 3 ( you see that there are a total of four 3's) is you need to know 3 exponent 3 3 * 3 *3 TIP: the exponent is the number of times the number gets multiplied. Im not familiar with the TI-34 but I think it should have this sign on it: ^ For example, if you need to know 3 exponent 4, just type in 3 ^ 4 and you should get an answer.. Hope this helps!

OpenStudy (anonymous):

Ok thanks sooooo much!! You are such a big help! I have fanned you! :)

OpenStudy (anonymous):

Haha, youre welcome! And thank you :D Im new here, so i'm happy I can help.. 

OpenStudy (anonymous):

Well, Welcome to Open Study!!